OpenStudy (sonyalee77):

Have vehicles gotten more fuel-efficient over the years? Between 1960 and 1999 the size and shape of automobiles in the United States has changed almost annually. The amount of fuel consumed by these vehicles has also changed. The following table describes the average fuel consumption per year per passenger car in gallons of gasoline. Year, t 1960 1970 1980 1990 1995 1997 1998 1999 Gallons Consumed Per Passenger Car 668 760 576 520 530 538 544 552 Determine the average rate of change, in gallons of fuel used per passenger car, from 1960 to 1970.

OpenStudy (sonyalee77):

@superdavesuper

OpenStudy (anonymous):

@inowalst @iambatman @OpenStudyRocks5* @YOLO_KING @satellite73 @triciaal @Tazmaniadevil any help on this question? D:

OpenStudy (inowalst):

Sorry, I dont know.

OpenStudy (yolo_king):

Have you tried making a scatterd chart with the given information?

OpenStudy (openstudyrocks5*):

Whats the Anwsers and what do you think it is?

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

@ParthKohli @ilovewolf @Destinymasha @Destinymasha @divu.mkr @Conqueror @bohotness @Zzero0 @undeadknight26 @pitamar anyone????

OpenStudy (openstudyrocks5*):

@iGreen

OpenStudy (anonymous):

no clue

OpenStudy (anonymous):

I'd say: $$ \frac{\Delta g}{\Delta t} = \frac{760 - 668}{1970 - 1960} = \frac{92}{10} = 9.2 $$

OpenStudy (anonymous):

Thank you!